# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ Abstract.
This paper is a complement of the modularity result of [${}^{+}$20] for the special case $U(1,1)$ not considered there. The main idea to embed a $U(1,1)$ Shimura curve to many $U(n-1,1)$ Shimura varieties for big $n$, and prove a precise pullback formula of the generating series of arithmetic divisors. Afterwards, we use the modularity result of [${}^{+}$20] together with existence of non-vanishing of classical theta series at any given point in the upper half plane to prove the modulartiy result on $U(1,1)$ Shimura curves.
@@
QH was partially supported by Columbia’s Faculty Research Allocation Program, YS is partially supported by Zhejiang University’s start-up fund, and TY was partially supported by UW-Madison Kellet Mid-career award
@@ 1. Introduction
where $r_{m}(n)$ is the number of ways to express $n$ as sum of $m$ integers. The modularity of $\theta_{m}$ implies in particular that if we know a few $r_{m}(n)$ for “small” $n$, then we know all $r_{m}(n)$. Modularity has played an important role in number theory and other fields, see for example [HZ76], [KM90], [Kud97b], [Zha21], [BWR15], and [${}^{+}$20].
In [${}^{+}$20], the authors proved the modularity of a generating series of special divisors on a compactified integral model of a Shimura variety associated to a unitary group (Krämer model for a unimodular lattice) of signature $(n-1,1)$ for $n>2$, and left the special case $n=2$ unresolved for technical reasons. This paper finishes the job for $n=2$ by the so-called embedding trick. We now describe the main results in a little more detail. For undefined terms, see Section 2 for definition.
Let $\boldsymbol{k}=\mathbb{Q}(\sqrt{-D})$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{\boldsymbol{k}}$ and odd discriminant $-D$. Let $\mathfrak{a}_{0}$ and $\mathfrak{a}$ be two unimodular Hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattices of signature $(1,0)$ and $(n-1,1)$ respectively. Let $\mathcal{S}^{*}$ be the associated integral model of the Shimura variety with toroidal compactification (see Section 2). For each $m>0$, Kudla and Rapoport ([KR14]) constructed divisors $\mathcal{Z}(m)$ in the open Shimura variety $\mathcal{S}\subset\mathcal{S}^{*}$ over $\mathcal{O}_{\boldsymbol{k}}$. There are two Green functions for $Z(m)=\mathcal{Z}(m)(\mathbb{C})$, $\mathrm{Gr}_{B}(m)$ constructed by J. Bruinier in his thesis ([Bru02], [BF04]), and $\mathrm{Gr}_{K}(m,v)$ constructed by S. Kudla with an extra parameter $v>0$ ([Kud97a]). By studying their behavior at the boundary, one obtains two arithmetic divisors ([BHY15], [How15], [ES18], see Sections 5 and 6)
@@
One also defines arithmetic divisors $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(0)$ and $\widehat{\mathcal{Z}}^{\mathrm{tot}}_{K}(0,v)$ in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$. Now we have two arithmetic theta functions with values in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$.
(1.2) $\widehat{\Theta}_{B}(\tau)$ $=\sum_{m\geq 0}\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m)q^{m}\in\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})[[q]],$
(1.3) $\widehat{\Theta}_{K}(\tau)$ $=\sum_{m\in\mathbb{Z}}\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(m,v)q^{m}\in\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})[[q,q^{-1}]].$
Ehlen and Sankran proved in [ES18, Theorem 1.4] that $\widehat{\Theta}_{B}(\tau)-\widehat{\Theta}_{K}(\tau)$ is a modular form for $\Gamma_{0}(D)$ of weight $n$, character $\chi_{-D}^{n}$ (where $\chi_{-D}(\cdot)$ is the Kronecker symbol $\left(\frac{-D}{\cdot}\right)$) and with values in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$ (see Section 7 for detail and some subtlety). In [${}^{+}$20], the authors proved the modularity of $\widehat{\Theta}_{B}(\tau)$ for $n\geq 3$, and thus the modularity of $\widehat{\Theta}_{K}(\tau)$ for $n\geq 3$. In this paper, we prove the modularity of $\widehat{\Theta}_{K}(\tau)$ for $n=2$, which in turn implies the modularity of $\widehat{\Theta}_{B}(\tau)$. We record it as
@@
We remark that the modularity of $\widehat{\Theta}_{B}$ could be used to prove a complement of the original Gross-Zagier formula (the case when every $p|N$ is ramified in $\boldsymbol{k}$, in comparison to that every $p|N$ is split in $\boldsymbol{k}$). Since $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$ is infinite dimensional, we refer to [ES18, Definition 4.11] for the precise meaning of the modularity. To prove the modularity of $\widehat{\Theta}_{K}(\tau)$ for $n=2$, we use the following embedding trick (see, for example, [Li23], [HMP20]). Let $\Lambda$ be a positive definite unimodular $\mathcal{O}_{\boldsymbol{k}}$-lattice of rank $m$. Replacing $\mathfrak{a}$ by $\mathfrak{a}\obot\Lambda$, we obtain a Shimura variety $\mathcal{S}^{\diamond,*}$ together with a canonical embedding:
(1.4) $\varphi_{\Lambda}:\mathcal{S}^{*}\rightarrow\mathcal{S}^{\diamond,*},(A_{0},A,...)\mapsto(A_{0},A\times(A_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda),...).$
@@
is a classical holomorphic modular form for $\Gamma_{0}(D)$ of weight $m$ and character $\chi_{-D}^{m}$
By an analogue of a result of Yingkun Li [Li23, Lemma 3.2], for every $\tau$, there is some $\Lambda$ with $\theta_{\Lambda}(\tau)\neq 0$. So $\widehat{\Theta}_{K}(\tau)$ is everywhere defined and the modularity of $\widehat{\Theta}_{K}(\tau)$ follows from the modularity result of [${}^{+}$20]. Now Theorem 1.1 for $\widehat{\Theta}_{B}(\tau)$ follows from the modularity for $\widehat{\Theta}_{K}(\tau)$ and the main result of [ES18]. Note that the analogue of Theorem 1.2 does not hold for $\widehat{\Theta}_{B}$.
The main work to prove Theorem 1.2 is to understand the pullback of special divisors on $\mathcal{S}^{\diamond,*}$ that intersect with $\mathcal{S}^{*}$ improperly and various subtleties about the boundary components. On the generic fiber, this pullback is controlled by the line bundle of modular form of weight $1$ (descended from the tautological line bundle), which can be regarded as an analogue of the adjunction formula. The line bundle of modular form of weight $1$ has two well-defined integral models: $\omega$ considered in [${}^{+}$20] and $\Omega$ considered in [How19]. Although $\omega$ is used in the definition of $\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(0,v)$ and $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(0)$ in [${}^{+}$20], it turns out that $\Omega$ is the one that controls the pullback of special divisors that intersect with $\mathcal{S}^{*}$ improperly. So it is important to figure out their precise relation. We show in Section 3 (Theorem 3.2):
@@
We believe that this relation is of independent interest. It is discovered by [How19] that the line bundle $\Omega$ controls the deformation theory of special divisors in the open Shimura variety $\mathcal{S}$. Another main technical point of this paper is to extend this observation to the toroidal compactification $\mathcal{S}^{*}$, see Theorem 4.4.
This paper is organized as follows. In Section 2, we review basics on integral model, toroidal compactification and special divisors which are needed in this paper. In Section 3, we prove the precise relation (1.6) between the two well-defined integral models of the line bundle of modular forms of weight $1$ (Theorem 3.2). In Section 4, we extend the controlling deformation property of $\Omega$ of Ben Howard to boundaries (Theorem 4.4). Section 5 is dedicated to proving Theorem 5.5: Theorem 1.2 without Green functions. In Section 6, we deal with pullbacks of Green functions and metrics on line bundles, and finish the proof of Theorem 1.2. In the first part of Section 7, we use Li’s embedding trick ([Li23]) to prove the modularity of $\widehat{\Theta}_{K}(\tau)$. In the second part of Section 7, we describe two slightly different Green functions for special divisors, one defined by Bruinier in [${}^{+}$20] and the other one defined by Ehlen and Sankaran in [ES18]. We show that $\widehat{\Theta}_{B}(\tau)$ and $\widehat{\Theta}_{ES}(\tau)$ are essentially the same (7.21), which will be used in the modularity of difference $\widehat{\Theta}_{K}(\tau)-\widehat{\Theta}_{B}(\tau)$. We remark that Bruinier’s Green functions are a little easier to define than those of Ehlen and Sankaran.
Acknowledgements: We thank Jan Bruinier, Ben Howard, Yingkun Li, Keerthi Madapusi Pera, and Sid Sankaran their helpful discussions during the preparation of this paper.
@@ 1.1. Notations
Let $\boldsymbol{k}=\mathbb{Q}(\sqrt{-D})$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{\boldsymbol{k}}$ and odd fundamental discriminant $-D$. Let $\chi_{-D}(\cdot)$ be the Kronecker symbol $\left(\frac{-D}{\cdot}\right)$. Fix a $\pi\in\mathcal{O}_{\boldsymbol{k}}$ such that $\mathcal{O}_{\boldsymbol{k}}=\mathbb{Z}+\pi\mathbb{Z}$. For any $\mathcal{O}_{\boldsymbol{k}}$-scheme $S$, define
@@ 2.1. Unitary Shimura varieties
In this section, we review the theory of unitary Shimura variety following [${}^{+}$20]. Let $W_{0}$ and $W$ be $\boldsymbol{k}$-vector spaces endowed with hermitian forms $H_{0}$ and $H$ of signatures $(1,0)$ and $(n-1,1)$, respectively. We always assume that $n\geqslant 2$. Abbreviate
@@
carries a hermitian form of signature $(n-1,1)$ defined by
@@
which is a unimodular hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattice of signature $(n-1,1)$. The group $G$ acts on $V$ in a natural way, defining an exact sequence
(2.6) $1\rightarrow\mathrm{Res}_{\boldsymbol{k}/\mathbb{Q}}\mathbb{G}_{m}\rightarrow G\rightarrow\mathrm{U}(V)\rightarrow 1$
@@
is the set of complex points of a smooth $\boldsymbol{k}$-stack of dimension $n-1$, denoted $\operatorname{Sh}(G,\mathcal{D})$.
@@ 2.2. Integral model
We recall the integral model defined by [${}^{+}$20] originated from the work of Krämer [Krä03]. For $(r,s)=(n-1,1)$ or $(n,0)$, let
@@
In particular, when $(r,s)=(n,0)$, the $\mathcal{F}_{A}$ in the condition $(4)$ above is simply $\operatorname{Lie}(A)$. In this case, $\mathcal{M}(n,0)\rightarrow\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$ is proper and smooth of relative dimension 00 by [How15, Proposition 2.1.2]. For $\mathcal{M}_{(n-1,1)}$, we have the following theorem.
@@
[${}^{+}$20, Theorem 2.3.3] The $\mathcal{O}_{\boldsymbol{k}}$-stack $\mathcal{M}_{(n-1,1)}$ is regular and flat with reduced fibers.
@@
[…]
@@
for every $\ell\neq\mathrm{char}(k(s))$. Here the hermitian form on the right hand side is defined as in (2.4). For the hermitian form on the left hand side, we define it similarly by replacing the symplectic forms (2.1) on $W_{0}$ and $W$ with the Weil pairings on the Tate modules $T_{\ell}A_{0,s}$ and $T_{\ell}A_{s}$ ($A_{s}$ being the pullback of the universal object $\mathbf{A}$ over $\mathcal{S}$ to $s$) induced by their polarizations. The generic fiber $\mathcal{S}\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}}\mathrm{Spec}\,\boldsymbol{k}$ is the Shimura variety $\operatorname{Sh}(G,\mathcal{D})$.
Now assume that $p\in\mathbb{Z}$ is a prime ramified in $\boldsymbol{k}$ (i.e. dividing $D$). Following [How19, Appendix A] we define the exceptional divisor $\mathrm{Exc}_{p}$ to be the locus of $\mathcal{S}\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k},p}}\mathrm{Spec}\,\mathbb{F}_{p}$ where a geometric point $s\in\mathcal{S}(\bar{\mathbb{F}}_{p})$ is in $\mathrm{Exc}_{p}$ if the action $\mathcal{O}_{\boldsymbol{k}}\rightarrow\operatorname{Lie}A_{s}$ factor through the reduction homomorphism $\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathbb{F}_{p}$. Then $\mathrm{Exc}_{p}$ can be given the structure of a reduced substack of $\mathcal{S}$ and is in fact a Cartier divisor consisting of disjoint unions of $\mathbb{P}^{n-1}$ over $\mathbb{F}_{p}$ ([${}^{+}$20, Theorem 2.3.4], [How19, Proposition A.2]). Finally, we define

### v2

@@ Abstract.
This paper is a complement of the modularity result of [BHK+20] for the special case $U(1,1)$ not considered there. The main idea is to embed a $U(1,1)$ Shimura curve to many U(n1,1)U(n-1,1) Shimura varieties for big $n$, and prove a precise pullback formula of the generating series of arithmetic divisors. Afterwards, we use the modularity result of [BHK+20] together with the existence of non-vanishing classical theta series at any given point in the upper half plane to prove the modulartiy result on $U(1,1)$ Shimura curves.
@@
QH was partially supported by Columbia’s Faculty Research Allocation Program, YS is partially supported by Zhejiang University’s start-up Fund and Qizhen Fund, and TY was partially supported by UW-Madison Kellet Mid-career award
@@ 1. Introduction
where $r_{m}(n)$ is the number of ways to express $n$ as sum of $m$ integers. The modularity of $\theta_{m}$ implies in particular that if we know a few $r_{m}(n)$ for “small” $n$, then we know all $r_{m}(n)$. Modularity has played an important role in number theory and other fields, see for example [HZ76], [KM90], [Kud97a], [Zha21], [BWR15], and [BHK+20].
In [BHK+20], the authors proved the modularity of a generating series of special divisors on a compactified integral model of a Shimura variety associated to a unitary group (Krämer model for a unimodular lattice) of signature (n1,1)(n-1,1) for $n>2$, and left the special case $n=2$ unresolved for technical reasons. This paper finishes the job for $n=2$ by the so-called embedding trick. We now describe the main results in a little more detail. For undefined terms, see Section 2 for definition.
Let $\boldsymbol{k}=\mathbb{Q}(\sqrt{-D})$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{\boldsymbol{k}}$ and odd discriminant D-D. Let $\mathfrak{a}_{0}$ and $\mathfrak{a}$ be two unimodular Hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattices of signature $(1,0)$ and (n1,1)(n-1,1) respectively. Let $\mathcal{S}^{*}$ be the associated integral model of the Shimura variety with toroidal compactification (see Section 2). For each $m>0$, Kudla and Rapoport ([KR14]) constructed divisors $\mathcal{Z}(m)$ in the open Shimura variety $\mathcal{S}\subset\mathcal{S}^{*}$ over $\mathcal{O}_{\boldsymbol{k}}$. There are two Green functions for $Z(m)=\mathcal{Z}(m)(\mathbb{C})$, $\mathrm{Gr}_{B}(m)$ constructed by J. Bruinier in his thesis ([Bru02], [BF04]), and $\mathrm{Gr}_{K}(m,v)$ constructed by S. Kudla with an extra parameter $v>0$ ([Kud97b]). By studying their behavior at the boundary, one obtains two arithmetic divisors ([BHY15], [How15], [ES18], see Sections 5 and 6)
@@
One also defines arithmetic divisors $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(0)$ and $\widehat{\mathcal{Z}}^{\mathrm{tot}}_{K}(0,v)$ in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$. As in [BHK+20], We define two formal $q$-expansions with values in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$:
(1.2) $\widehat{\Theta}_{B}(\tau)$ $=\sum_{m\geq 0}\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m)q^{m},\hbox{ and }$
(1.3) $\widehat{\Theta}_{K}(\tau)$ $=\sum_{m\in\mathbb{Z}}\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(m,v)q^{m}.$
They are usually called Bruinier and Kudla arithmetic theta functions respectively. Ehlen and Sankran proved in [ES18, Theorem 1.4] that $\widehat{\Theta}_{B}(\tau)-\widehat{\Theta}_{K}(\tau)$ is a modular form for $\Gamma_{0}(D)$ of weight $n$, character $\chi_{-D}^{n}$ (where $\chi_{-D}(\cdot)$ is the Kronecker symbol $\left(\frac{-D}{\cdot}\right)$) and with values in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$ (see Section 7 for detail and some subtlety). In [BHK+20], the authors proved the modularity of $\widehat{\Theta}_{B}(\tau)$ for $n\geq 3$, and thus the modularity of $\widehat{\Theta}_{K}(\tau)$ for $n\geq 3$. In this paper, we prove the modularity of $\widehat{\Theta}_{K}(\tau)$ for $n=2$, which in turn implies the modularity of $\widehat{\Theta}_{B}(\tau)$. We record it as
@@
Remark 1.2.
As $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$ is infinitely dimensional, the modularity needs some explanation. Since $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m)$ is independent of $\tau$, $\widehat{\Theta}_{B}(\tau)$ being modular means one of the two equivalent conditions:
(1) For every linear functional $f$ on $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$, $f(\widehat{\Theta}_{B}(\tau))=\sum f(\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m))q^{m}$ is modular.
(2) There are finitely many arithmetic divisors $\widehat{Z}_{i}$ such that
$\widehat{\Theta}_{B}(\tau)=\sum_{i}f_{i}(\tau)\widehat{\mathcal{Z}}_{i}$
and $f_{i}(\tau)$ are usual scalar holomorphic modular forms.
However, since $\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(m,v)$ depends on $v=\hbox{Im}(\tau)$, the above definition does not work and the modularity is more subtle. We refer to [ES18, Definition 4.1] for detail. Roughly speaking, it means that we can write
$\widehat{\Theta}_{K}(\tau)=\sum_{i}f_{i}(\tau)\widehat{\mathcal{Z}}_{i}+(0,f(\tau,z))$
with $f_{i}$ and $\widehat{\mathcal{Z}}_{i}$ as above, and that $f(\tau,z)$ is a smooth scalar modular form for every $z\in S$ plus some extra technical conditions.
We remark that the modularity of $\widehat{\Theta}_{B}$ could be used to prove a complement of the original Gross-Zagier formula (the case when every $p|N$ is ramified in $\boldsymbol{k}$, in comparison to that every $p|N$ is split in $\boldsymbol{k}$). To prove the modularity of $\widehat{\Theta}_{K}(\tau)$ for $n=2$, we use the following embedding trick (see, for example, [Li23], [HMP20]). Let $\Lambda$ be a positive definite unimodular $\mathcal{O}_{\boldsymbol{k}}$-lattice of rank $m$. Replacing $\mathfrak{a}$ by $\mathfrak{a}\obot\Lambda$, we obtain a Shimura variety $\mathcal{S}^{\diamond,*}$ together with a canonical morphism:
[…]
@@
is a classical holomorphic modular form for $\Gamma_{0}(D)$ of weight $m$ and character $\chi_{-D}^{m}$ .
By an analogue of a result of Yingkun Li [Li23, Lemma 3.2], for every $\tau$, there is some $\Lambda$ with $\theta_{\Lambda}(\tau)\neq 0$. So $\widehat{\Theta}_{K}(\tau)$ is everywhere defined and the modularity of $\widehat{\Theta}_{K}(\tau)$ follows from the modularity result of [BHK+20]. Now Theorem 1.1 for $\widehat{\Theta}_{B}(\tau)$ follows from the modularity for $\widehat{\Theta}_{K}(\tau)$ and the main result of [ES18]. Note that the analogue of Theorem 1.3 does not hold for $\widehat{\Theta}_{B}$.
The main work to prove Theorem 1.3 is to understand the pullback of special divisors on $\mathcal{S}^{\diamond,*}$ that intersect with $\mathcal{S}^{*}$ improperly and various subtleties about the boundary components. On the generic fiber, this pullback is controlled by the line bundle of modular form of weight $1$ (descended from the tautological line bundle), which can be regarded as an analogue of the adjunction formula. The line bundle of modular form of weight $1$ has two well-defined integral models: $\omega$ considered in [BHK+20] and $\Omega$ considered in [How19]. Although $\omega$ is used in the definition of $\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(0,v)$ and $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(0)$ in [BHK+20], it turns out that $\Omega$ is the one that controls the pullback of special divisors that intersect with $\mathcal{S}^{*}$ improperly. So it is important to figure out their precise relation. We show in Section 3 (Theorem 3.2):
@@
We believe that this relation is of independent interest too. It is discovered by [How19] that the line bundle $\Omega$ controls the deformation theory of special divisors in the open Shimura variety $\mathcal{S}$. Another main technical point of this paper is to extend this observation to the toroidal compactification $\mathcal{S}^{*}$, see Proposition 4.8. In order to do this, we extend the definition of $\mathcal{Z}(m)$ as a stack by [KR14] to the boundary, see Definition 4.5.
This paper is organized as follows. In Section 2, we review basics on integral model, toroidal compactification and special divisors which are needed in this paper. In Section 3, we prove the precise relation (1.6) between the two well-defined integral models of the line bundle of modular forms of weight $1$ (Theorem 3.2). In Section 4, we extend the controlling deformation property of $\Omega$ of Ben Howard to boundaries (Proposition 4.8). Section 5 is dedicated to proving Theorem 5.8: Theorem 1.3 without Green functions. In Section 6, we deal with pullbacks of Green functions and metrics on line bundles, and finish the proof of Theorem 1.3. In the first part of Section 7, we use Li’s embedding trick ([Li23]) to prove the modularity of $\widehat{\Theta}_{K}(\tau)$. In the second part of Section 7, we describe two slightly different Green functions for special divisors, one defined by Bruinier in [BHK+20] and the other one defined by Ehlen and Sankaran in [ES18]. We show that $\widehat{\Theta}_{B}(\tau)$ and $\widehat{\Theta}_{ES}(\tau)$ are essentially the same (7.20), which will be used in the modularity of difference $\widehat{\Theta}_{K}(\tau)-\widehat{\Theta}_{B}(\tau)$. We remark that Bruinier’s Green functions are a little easier to define than those of Ehlen and Sankaran.
Acknowledgment: We thank Jan Bruinier, Ben Howard, Yingkun Li, Keerthi Madapusi Pera, and Sid Sankaran their helpful discussions during the preparation of this paper. We thank the anonymous referee for carefully reading the paper and for their valuable suggestion and comments, which makes the paper better written.
@@ 1.1. Notations
Let $\boldsymbol{k}=\mathbb{Q}(\sqrt{-D})$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{\boldsymbol{k}}$ and odd fundamental discriminant D-D. Let $\chi_{-D}(\cdot)$ be the Kronecker symbol $\left(\frac{-D}{\cdot}\right)$. Fix a $\pi\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}=\mathbb{Z}+\pi\mathbb{Z}$. We also set $\delta=\sqrt{-D}$. For any $\mathcal{O}_{\boldsymbol{k}}$-scheme $S$, define
@@ 2.1. Unitary Shimura varieties
In this section, we review the theory of unitary Shimura variety following [BHK+20]. Let $W_{0}$ and $W$ be $\boldsymbol{k}$-vector spaces endowed with hermitian forms $H_{0}$ and $H$ of signatures $(1,0)$ and (n1,1)(n-1,1), respectively. We always assume that $n\geqslant 2$. Abbreviate
@@
carries a hermitian form of signature (n1,1)(n-1,1) defined by
@@
which is a unimodular hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattice of signature (n1,1)(n-1,1). The group $G$ acts on $V$ in a natural way, defining an exact sequence
(2.6) $1\rightarrow\mathrm{Res}_{\boldsymbol{k}/\mathbb{Q}}\mathbb{G}_{m}\rightarrow G\rightarrow\mathrm{U}(V)\rightarrow 1.$
@@
is the set of complex points of a smooth $\boldsymbol{k}$-stack of dimension n1n-1, denoted by $\operatorname{Sh}(G,\mathcal{D})$.
@@ 2.2. Integral model
We recall the integral model defined by [BHK+20] originating in the work of Krämer [Krä03]. For (r,s)=(n1,1)(r,s)=(n-1,1) or $(n,0)$, let
@@
In particular, when $(r,s)=(n,0)$, the $\mathcal{F}_{A}$ in the condition $(4)$ above is simply $\operatorname{Lie}(A)$. In this case, $\mathcal{M}(n,0)\rightarrow\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$ is proper and smooth of relative dimension 0 by [How15, Proposition 2.1.2]. For $\mathcal{M}_{(n-1,1)}$, we have the following theorem.
@@
[BHK+20, Theorem 2.3.3] The $\mathcal{O}_{\boldsymbol{k}}$-stack $\mathcal{M}_{(n-1,1)}$ is regular and flat with reduced fibers.
@@
to be the open and closed substack such that $\mathcal{S}(S)$ is the groupoid of tuples
@@
for every finite prime $\ell\neq\mathrm{char}(k(s))$. Here the hermitian form on the right hand side is defined as in (2.4). For the hermitian form on the left hand side, we define it similarly by replacing the symplectic forms (2.1) on $W_{0}$ and $W$ with the Weil pairings on the Tate modules $T_{\ell}A_{0,s}$ and $T_{\ell}A_{s}$ ($A_{s}$ being the pullback of the universal object $\mathbf{A}$ over $\mathcal{S}$ to $s$) induced by their polarizations. The generic fiber $\mathcal{S}\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}}\mathrm{Spec}\,\boldsymbol{k}$ is the Shimura variety $\operatorname{Sh}(G,\mathcal{D})$.
Now assume that $p\in\mathbb{Z}$ is a prime ramified in $\boldsymbol{k}$ (i.e. dividing $D$). Following [How19, Appendix A] we define the exceptional divisor $\mathrm{Exc}_{p}$ to be the locus of $\mathcal{S}\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k},p}}\mathrm{Spec}\,\mathbb{F}_{p}$ where a geometric point $s\in\mathcal{S}(\bar{\mathbb{F}}_{p})$ is in $\mathrm{Exc}_{p}$ if the action $\mathcal{O}_{\boldsymbol{k}}\rightarrow\operatorname{Lie}A_{s}$ factor through the reduction homomorphism $\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathbb{F}_{p}$. Then $\mathrm{Exc}_{p}$ can be given the structure of a reduced substack of $\mathcal{S}$ and is in fact a Cartier divisor consisting of disjoint unions of $\mathbb{P}^{n-1}$ over $\mathbb{F}_{p}$ ([BHK+20, Theorem 2.3.4], [How19, Proposition A.2]). Finally, we define